OpenStudy (anonymous):

How do you find all the real solutions to x^3-13x-12=0?

OpenStudy (anonymous):

wait one moment

OpenStudy (anonymous):

\[x \approx4 ; x \approx-3 ; x \approx-1\]

OpenStudy (anonymous):

I was a bit complicated

OpenStudy (anonymous):

was rectified and that is the correct answer

OpenStudy (anonymous):

What about x^4-14x^2+45=0?

OpenStudy (anonymous):

wait one moment

OpenStudy (dumbcow):

factor (x^2 -5)(x^2 -9)=0 x^2 = 5 x^2 = 9

OpenStudy (anonymous):

x=3 x=-3 x=square root(5) x=-square root(5)

OpenStudy (anonymous):

\[\sqrt{5}\approx 2.23 ; -\sqrt{5}\approx - 2.23\]

OpenStudy (anonymous):

Could you still factor it if the 14x^2 was 14x^3(in the same problem)?

OpenStudy (anonymous):

wait one moment

OpenStudy (anonymous):

is very complicated

OpenStudy (anonymous):

wait one moment

OpenStudy (anonymous):

Don't worry about it then. I was just wondering if they could be only one power different or if they could also be two powers away from each other.

OpenStudy (anonymous):

solve this equation is very long process and I think I'm doing that I have out there in 10%

OpenStudy (anonymous):

expected and it started to let me finish lol

OpenStudy (anonymous):

Can you show me the steps to the x^4-7x^3=0 problem? i think I know how to get 7 but I don't know how you got sero as an answer.

OpenStudy (anonymous):

only by deduction. in looking for possible solutions when the right side of the expression as a result of 0 in this case is very easy to see that 7 is the only number other than 0 that satisfies this condition and so that the other three answers are 0 and which is a 4-degree equation must give 4 answers

OpenStudy (anonymous):

Okay, so you get x^3(x-7)=0 and solve the two pieces for the answers?

OpenStudy (anonymous):

( x^3 ) (x-7)= 0 ??????

OpenStudy (anonymous):

I think that is how I was taught it. Then you get x=7 and x^3 =0